OpenStudy (anonymous):

Another one: Moivre theorem: (sqrt2cis(pi/9)^6... now what am I doing wrong?

OpenStudy (anonymous):

\[\sqrt{2}(\cos(\frac{ \pi }{9})n+isin(\frac{ \pi }{ 9 }))^{6} = 6(\cos \frac{ 2\pi }{ 3 }+isin \frac{ 2\pi }{3 })\]

zepdrix (zepdrix):

six square roots of 2 should give us 2x2x2, yes? Check your multiplication there :)

OpenStudy (anonymous):

6(-1/2 + sqrt3 /2 i) = -3+3sqrt3i

OpenStudy (anonymous):

oooo so i put 6 instead of 8

zepdrix (zepdrix):

ya that's the only mistake i see :D

OpenStudy (anonymous):

omgosh my mistakes are soooo dumb... thank u

zepdrix (zepdrix):

hehe :3

OpenStudy (anonymous):

thanks!